OpenStudy (anonymous):

Find the arc length of the curve x = cos(3t), y = sin(3t) where 0 ≤ t ≤ π/2.

OpenStudy (anonymous):

dx/dt = -3sin(3t) and dy/dt= 3cos(3t)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} \sqrt{(-3\sin(3t))^2 +(3\cos(3t))^2}\]

OpenStudy (anonymous):

i can't simplify this to find its integration

OpenStudy (anonymous):

would the square root be simplified to just 3?

OpenStudy (anonymous):